# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 7. Thickened $K_{4,m}$’s
This section deals with the thickened $K_{4,m}$ outcome of Theorem 1; see 7.8.
@@
Let $Z\subseteq V(G)$ and $K$ a component of $G-Z$. When $Z$ and $K$ are clear from context, we write $(U^{\prime}_{K},W^{\prime}_{K}):=(V(K)\cup Z,V(G-K))$. Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. For each component $K$ of $G-Z$, let $(U_{K},W_{K})$ denote the left-right-shift of $(U^{\prime}_{K},W^{\prime}_{K})$.
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$, such that $G-Z$ has $\geqslant 4$ components. Let $K$ be a component of $G-Z$. Then $(U_{K},W_{K})$ equals the left-shift of $(U^{\prime}_{K},W^{\prime}_{K})$. In particular, $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ and $W_{K}=W^{\prime}_{K}$.
@@
Let $(A,B)$ denote the left-shift of $(U^{\prime}_{K},W^{\prime}_{K})$. Since every vertex in $S(U^{\prime}_{K},W^{\prime}_{K})$ has $\geqslant 2$ neighbours in $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}$, and since $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}\subseteq B\smallsetminus A$ by 5.2, it follows that every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$. Hence $(A,B)=(U_{K},W_{K})$ and $W^{\prime}_{K}=B=W_{K}$. ∎
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Let $K$ be a component of $G-Z$. Then the following assertions are equivalent:
@@
    the left-right-shift of $(U^{\prime}_{K},W^{\prime}_{K})$ is a totally-nested tetra-separation.
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). If $K$ has only one vertex, then $|U_{K}\smallsetminus W_{K}|=1$, so $(U^{\prime}_{K},W^{\prime}_{K})$ fails to be a tetra-separation by 5.10.
(i) $\Rightarrow$ (ii). We abbreviate $(U,W):=(U_{K},W_{K})$ and $(U^{\prime},W^{\prime}):=(U^{\prime}_{K},W^{\prime}_{K})$. Since $K$ has $\geqslant 2$ vertices, $(U,W)$ is a tetra-separation by 5.10. To show that $(U,W)$ is totally-nested, by 6.16 it suffices to show that $(U,W)$ is externally 5-connected.
Half-connected: We have $G[U\smallsetminus W]=K$ by Lemma 7.1.
To check the remaining aspects of external 5-connectivity, for each $z\in Z$ we denote by $z^{*}$ the element of $S(U,W)$ that equal or incident to $z$. By Lemma 7.1, $z^{*}$ is either equal to $z$ or $z^{*}$ is an edge between $K$ and $z$. Denote by $z^{U}$ the vertex in $\hat{z}\cap U$. We can use the three components of $G-Z$ besides $K$ to find paths:
@@
3-linked: follows from ($\dagger$)Section 7.
0-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $W\smallsetminus U$ by ($\dagger$)Section 7. We claim that there are two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $K$. Let $\{\pi^{\prime}_{1},\pi^{\prime}_{2}\}$ denote the bipartition of $Z$ that corresponds to $\{\pi_{1},\pi_{2}\}$. For both $i$, let $\pi_{i}^{K}$ denote the set of neighbours of $\pi^{\prime}_{i}$ in $K$. Then $|\pi_{i}^{K}|\geqslant 2$, since otherwise $\pi_{i}^{K}\cup\pi^{\prime}_{3-i}$ is a 3-separator of $G$ as $|K|\geqslant 2$, contradicting 4-connectivity. So it suffices to find two disjoint $\pi_{1}^{K}$$\pi_{2}^{K}$ paths in $K$. If these are missing, then by Menger’s theorem $\pi_{1}^{K}$ and $\pi_{2}^{K}$ are separated by a single vertex $v$, and then $\pi^{\prime}_{1}\cup\{v\}$ is a 3-separator of $G$, contradicting 4-connectivity.
1-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Assume that $\pi_{1}$ is weird, say. Then $\pi_{1}$ is $U$-weird since each vertex in $Z$ has $\geqslant 3$ neighbours in $W\smallsetminus U$. By ($\dagger$)Section 7, there exist three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$. Hence $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(U,W)$.
@@
We write $)𝜎𝑍{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\sigma(Z)}}$ for the star $\{\,(U_{K},W_{K}):K$ is a component of $G-Z$ with $|K|\geqslant 2\,\}$.
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Then $\sigma(Z)$ is a star of totally-nested tetra-separations.
@@
By Lemma 7.2, every element in $\sigma(Z)$ is a totally-nested tetra-separation. Let $K\neq L$ be components of $G-Z$, each of size $\geqslant 2$. By Lemma 7.1, $(U_{K},W_{K})\leqslant(U_{K}^{\prime},W_{K}^{\prime})$ and $(U_{L},W_{L})\leqslant(U_{L}^{\prime},W_{L}^{\prime})$. With $(U_{K}^{\prime},W_{K}^{\prime})\leqslant(W_{L}^{\prime},U_{L}^{\prime})$, it follows that $(U_{K},W_{K})\leqslant(W_{L},U_{L})$. So, $\sigma(Z)$ is a star. ∎
@@
Let $(A,B)$ be a 4-separation in a 4-connected graph $G$ such that $|A\smallsetminus B|\geqslant 2$ and every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$. Let $(A^{\prime\prime},B^{\prime\prime})$ be the left-right-shift of $(A,B)$. Then no tetra-separation $(C,D)$ of $G$ satisfies $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$.
@@
Suppose for a contradiction that there is a tetra-separation $(C,D)$ with $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$. By 5.10, $(A^{\prime\prime},B^{\prime\prime})$ is a tetra-separation. By Lemma 7.1, we have $B=B^{\prime\prime}$. So both inclusions $B^{\prime\prime}\supseteq D\supseteq B$ are actually equalities. We therefore have $A^{\prime\prime}\subsetneq C\subseteq A$. Let $v\in C\smallsetminus A^{\prime\prime}$. Then $v$ is a vertex in $S(A,B)$ with at most one neighbour in $A\smallsetminus B$. But since $D=B$ we have $C\smallsetminus D\subseteq A\smallsetminus B$, so $v$ also has at most one neighbour in $C\smallsetminus D$, contradicting that $(C,D)$ is a tetra-separation. ∎
@@
By assumption, both $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ have two components. In particular, $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Then by Lemma 5.4$(A,B)$ satisfies the matching-condition. Let $v$ be a vertex in $S(A,B)$. By 4-connectivity of $G$ and by Lemma 6.1, $v$ has a neighbour in every component of $G-S(A,B)$. In particular, $v$ satisfies the degree-condition. Hence, $(A,B)$ is a tetra-separation. ∎
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that $G-Z$ has $\geqslant 4$ components. Let $(E,F)$ be a totally-nested tetra-separation of $G$. Then there exists a component $K$ of $G-Z$ with $|K|\geqslant 2$ such that $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ or $(F,E)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$.
@@
We denote by $\mathcal{K}$ the set of components of $G-Z$. First, we show that there is a component $K\in\mathcal{K}$ such that either $E\smallsetminus Z\subseteq V(K)$ or $F\smallsetminus Z\subseteq V(K)$. Assume otherwise. Then we can bipartition $\mathcal{K}$ into two sets $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ such that for both $i\in\{1,2\}$, $|\mathcal{K}_{i}|\geqslant 2$ and $\bigcup_{K\in\mathcal{K}_{i}}K$ intersects both $E$ and $F$. We define $A_{i}:=Z\cup\bigcup_{K\in\mathcal{K}_{i}}V(K)$. Then $(A_{1},A_{2})$ is a tetra-separation by Lemma 7.5. Now $E\nsubseteq A_{i}$ and $F\nsubseteq A_{i}$ for both $i\in\{1,2\}$ implies that $(A_{1},A_{2})$ crosses $(E,F)$, contradicting that $(E,F)$ is totally-nested.
Hence we have $E\subseteq V(K)\cup Z$, say, for a component $K\in\mathcal{K}$. Since every vertex in $Z$ has $\geqslant 2$ neighbours outside $V(K)\cup Z$, it follows that $(U^{\prime}_{K},W^{\prime}_{K})$ equals its right-shift. Hence $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma 5.7. Since $(E,F)$ is a tetra-separation, we have $|E\smallsetminus F|\geqslant 2$ by Lemma 5.5. Therefore, $|K|\geqslant 2$. ∎
@@
Let $G$ be a 4-connected graph, and let $N$ denote the set of totally-nested tetra-separations of $G$. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross so that the vertex-centre $Z$ has size four. Let $m$ denote the number of trivial components of $G-Z$. Then:
@@
    If $\sigma(Z)\neq\emptyset$, then $G-Z$ has a component with $\geqslant 2$ vertices, and the torso of $\sigma(Z)$ is a thickened $K_{4,m}$ where $Z$ is the left side.
@@
    If $\sigma(Z)=\emptyset$, then every component of $G-Z$ is trivial, and $G=\tau$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$.
@@
(i). By Corollary 7.3, $\sigma(Z)$ is a star in $N$. To show that $\sigma(Z)$ is splitting, assume for a contradiction that it is interlaced by a totally-nested tetra-separation $(E,F)$. By Lemma 7.6, there is a component $K$ of $G-Z$ with $|K|\geqslant 2$ and $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$, say. Since $K$ has at least two vertices, we have $(U_{K},W_{K})\in\sigma(Z)$. Note that $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma 7.1, which implies $(U_{K},W_{K})\not<(F,E)$. By Lemma 7.4, we cannot have $(U_{K},W_{K})<(E,F)$. Since $(U_{K},W_{K})\in N$, we therefore have $(E,F)\leqslant(U_{K},W_{K})$ or $(F,E)\leqslant(U_{K},W_{K})$, contradicting that $(E,F)$ interlaces $\sigma(Z)$.
(ii). We denote by $\sigma^{\prime}(Z)$ the star $\{\,(U^{\prime}_{K},W^{\prime}_{K}):K$ is a component of $G-Z$ with $|K|\geqslant 2\,\}$. Both $(A,B)$ and $(C,D)$ interlace $\sigma^{\prime}(Z)$, hence they also interlace $\sigma(Z)$ by Lemma 7.1.
(iii) and (iv) are straightforward. ∎
@@
(ii) $\Rightarrow$ (i). Let $X$ be a sprinkled side of $G$. Let $\{P_{1},\ldots,P_{4}\}$ be a partition of the other side $V(G-X)$. Let $(A,B)$ be the mixed-4-separation $(P_{1}\cup P_{2}\cup X,X\cup P_{3}\cup P_{4})$ and let $(C,D)$ be the mixed-4-separation $(P_{1}\cup P_{3}\cup X,X\cup P_{2}\cup P_{4})$. Then $(A,B)$ and $(C,D)$ are tetra-separations with separators equal to $X$, and they cross.
(i) $\Rightarrow$ (ii) follows from 7.7 (iii) and (iv). ∎

### v2

@@ 7. Thickened $K_{4,m}$’s
This section deals with the thickened $K_{4,m}$ outcome of Theorem˜1; see ˜7.8.
@@
Let $Z\subseteq V(G)$ and $K$ a component of GZG-Z. When $Z$ and $K$ are clear from context, we write $(U^{\prime}_{K},W^{\prime}_{K}):=(V(K)\cup Z,V(G-K))$. Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components. For each component $K$ of GZG-Z, let $(U_{K},W_{K})$ denote the left-right-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$.
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$, such that GZG-Z has $\geqslant 4$ components. Let $K$ be a component of GZG-Z. Then $(U_{K},W_{K})$ equals the left-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$. In particular, $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ and $W_{K}=W^{\prime}_{K}$.
@@
Let $(A,B)$ denote the left-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$. Since every vertex in $S(U^{\prime}_{K},W^{\prime}_{K})$ has $\geqslant 2$ neighbours in $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}$, and since $W^{\prime}_{K}\smallsetminus U^{\prime}_{K}\subseteq B\smallsetminus A$ by ˜5.2, it follows that every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$. Hence $(A,B)=(U_{K},W_{K})$ and $W^{\prime}_{K}=B=W_{K}$. ∎
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components. Let $K$ be a component of GZG-Z. Then the following assertions are equivalent:
@@
    the left-right-reduction of $(U^{\prime}_{K},W^{\prime}_{K})$ is a totally-nested tetra-separation.
@@
$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii). If $K$ has only one vertex, then $|U_{K}\smallsetminus W_{K}|=1$, so $(U^{\prime}_{K},W^{\prime}_{K})$ fails to be a tetra-separation by ˜5.10.
˜(i) $\Rightarrow$ ˜(ii). We abbreviate $(U,W):=(U_{K},W_{K})$ and $(U^{\prime},W^{\prime}):=(U^{\prime}_{K},W^{\prime}_{K})$. Since $K$ has $\geqslant 2$ vertices, $(U,W)$ is a tetra-separation by ˜5.10. To show that $(U,W)$ is totally-nested, by ˜6.16 it suffices to show that $(U,W)$ is externally 5-connected.
Half-connected: We have $G[U\smallsetminus W]=K$ by Lemma˜7.1.
To check the remaining aspects of external 5-connectivity, for each $z\in Z$ we denote by $z^{*}$ the element of $S(U,W)$ that equal or incident to $z$. By Lemma˜7.1, $z^{*}$ is either equal to $z$ or $z^{*}$ is an edge between $K$ and $z$. Denote by $z^{U}$ the vertex in $\hat{z}\cap U$. We can use the three components of GZG-Z besides $K$ to find paths:
@@
3-linked: follows from ˜($\dagger$)Section˜7.
0-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. We find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $W\smallsetminus U$ by ˜($\dagger$)Section˜7. We claim that there are two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths through $K$. Let $\{\pi^{\prime}_{1},\pi^{\prime}_{2}\}$ denote the bipartition of $Z$ that corresponds to $\{\pi_{1},\pi_{2}\}$. For both $i$, let $\pi_{i}^{K}$ denote the set of neighbours of $\pi^{\prime}_{i}$ in $K$. Then $|\pi_{i}^{K}|\geqslant 2$, since otherwise $\pi_{i}^{K}\cup\pi^{\prime}_{3-i}$ is a 3-separator of $G$ as $|K|\geqslant 2$, contradicting 4-connectivity. So it suffices to find two disjoint $\pi_{1}^{K}$$\pi_{2}^{K}$ paths in $K$. If these are missing, then by Menger’s theorem $\pi_{1}^{K}$ and $\pi_{2}^{K}$ are separated by a single vertex $v$, and then $\pi^{\prime}_{1}\cup\{v\}$ is a 3-separator of $G$, contradicting 4-connectivity.
1-potter-linked: Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Assume that $\pi_{1}$ is weird, say. Then $\pi_{1}$ is $U$-weird since each vertex in $Z$ has $\geqslant 3$ neighbours in $W\smallsetminus U$. By ˜($\dagger$)Section˜7, there exist three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$. Hence $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked with respect to $(U,W)$.
@@
[…]
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components. Then $\sigma(Z)$ is a star of totally-nested tetra-separations.
@@
By Lemma˜7.2, every element in $\sigma(Z)$ is a totally-nested tetra-separation. Let $K\neq L$ be components of GZG-Z, each of size $\geqslant 2$. By Lemma˜7.1, $(U_{K},W_{K})\leqslant(U_{K}^{\prime},W_{K}^{\prime})$ and $(U_{L},W_{L})\leqslant(U_{L}^{\prime},W_{L}^{\prime})$. With $(U_{K}^{\prime},W_{K}^{\prime})\leqslant(W_{L}^{\prime},U_{L}^{\prime})$, it follows that $(U_{K},W_{K})\leqslant(W_{L},U_{L})$. So, $\sigma(Z)$ is a star. ∎
@@
Let $(A,B)$ be a 4-separation in a 4-connected graph $G$ such that $|A\smallsetminus B|\geqslant 2$ and every vertex in $S(A,B)$ has $\geqslant 2$ neighbours in $B\smallsetminus A$. Let $(A^{\prime\prime},B^{\prime\prime})$ be the left-right-reduction of $(A,B)$. Then no tetra-separation $(C,D)$ of $G$ satisfies $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$.
@@
Suppose for a contradiction that there is a tetra-separation $(C,D)$ with $(A^{\prime\prime},B^{\prime\prime})<(C,D)\leqslant(A,B)$. By ˜5.10, $(A^{\prime\prime},B^{\prime\prime})$ is a tetra-separation. By Lemma˜7.1, we have $B=B^{\prime\prime}$. So both inclusions $B^{\prime\prime}\supseteq D\supseteq B$ are actually equalities. We therefore have $A^{\prime\prime}\subsetneq C\subseteq A$. Let $v\in C\smallsetminus A^{\prime\prime}$. Then $v$ is a vertex in $S(A,B)$ with at most one neighbour in $A\smallsetminus B$. But since $D=B$ we have $C\smallsetminus D\subseteq A\smallsetminus B$, so $v$ also has at most one neighbour in $C\smallsetminus D$, contradicting that $(C,D)$ is a tetra-separation. ∎
@@
By assumption, both $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ have two components. In particular, $|A\smallsetminus B|\geqslant 2$ and $|B\smallsetminus A|\geqslant 2$. Then by Lemma˜5.4(A,B)(A,B) satisfies the matching-condition. Let $v$ be a vertex in $S(A,B)$. By 4-connectivity of $G$ and by Lemma˜6.1, $v$ has a neighbour in every component of GS(A,B)G-S(A,B). In particular, $v$ satisfies the degree-condition. Hence, $(A,B)$ is a tetra-separation. ∎
@@
Let $Z$ be a set of four vertices in a 4-connected graph $G$ such that GZG-Z has $\geqslant 4$ components. Let $(E,F)$ be a totally-nested tetra-separation of $G$. Then there exists a component $K$ of GZG-Z with $|K|\geqslant 2$ such that $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ or $(F,E)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$.
@@
We denote by $\mathcal{K}$ the set of components of GZG-Z. First, we show that there is a component $K\in\mathcal{K}$ such that either $E\smallsetminus Z\subseteq V(K)$ or $F\smallsetminus Z\subseteq V(K)$. Assume otherwise. Then we can bipartition $\mathcal{K}$ into two sets $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ such that for both $i\in\{1,2\}$, $|\mathcal{K}_{i}|\geqslant 2$ and $\bigcup_{K\in\mathcal{K}_{i}}K$ intersects both $E$ and $F$. We define $A_{i}:=Z\cup\bigcup_{K\in\mathcal{K}_{i}}V(K)$. Then $(A_{1},A_{2})$ is a tetra-separation by Lemma˜7.5. Now $E\nsubseteq A_{i}$ and $F\nsubseteq A_{i}$ for both $i\in\{1,2\}$ implies that $(A_{1},A_{2})$ crosses $(E,F)$, contradicting that $(E,F)$ is totally-nested.
Hence we have $E\subseteq V(K)\cup Z$, say, for a component $K\in\mathcal{K}$. Since every vertex in $Z$ has $\geqslant 2$ neighbours outside $V(K)\cup Z$, it follows that $(U^{\prime}_{K},W^{\prime}_{K})$ equals its right-reduction. Hence $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma˜5.7. Since $(E,F)$ is a tetra-separation, we have $|E\smallsetminus F|\geqslant 2$ by Lemma˜$5.$ Therefore, $|K|\geqslant 2$. ∎
@@
Let $G$ be a 4-connected graph, and let $N$ denote the set of totally-nested tetra-separations of $G$. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross so that the vertex-centre $Z$ has size four. Let $m$ denote the number of trivial components of GZG-Z. Then:
@@
    If $\sigma(Z)\neq\emptyset$, then GZG-Z has a component with $\geqslant 2$ vertices, and the torso of $\sigma(Z)$ is a thickened $K_{4,m}$ where $Z$ is the left side.
@@
    If $\sigma(Z)=\emptyset$, then every component of GZG-Z is trivial, and $G=\tau$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$.
@@
˜(i). By Corollary˜7.3, $\sigma(Z)$ is a star in $N$. To show that $\sigma(Z)$ is splitting, assume for a contradiction that it is interlaced by a totally-nested tetra-separation $(E,F)$. By Lemma˜7.6, there is a component $K$ of GZG-Z with $|K|\geqslant 2$ and $(E,F)\leqslant(U^{\prime}_{K},W^{\prime}_{K})$, say. Since $K$ has at least two vertices, we have $(U_{K},W_{K})\in\sigma(Z)$. Note that $(U_{K},W_{K})\leqslant(U^{\prime}_{K},W^{\prime}_{K})$ by Lemma˜7.1, which implies $(U_{K},W_{K})\not<(F,E)$. By Lemma˜7.4, we cannot have $(U_{K},W_{K})<(E,F)$. Since $(U_{K},W_{K})\in N$, we therefore have $(E,F)\leqslant(U_{K},W_{K})$ or $(F,E)\leqslant(U_{K},W_{K})$, contradicting that $(E,F)$ interlaces $\sigma(Z)$.
˜(ii). We denote by $\sigma^{\prime}(Z)$ the star $\{\,(U^{\prime}_{K},W^{\prime}_{K}):K$ is a component of GZG-Z with $|K|\geqslant 2\,\}$. Both $(A,B)$ and $(C,D)$ interlace $\sigma^{\prime}(Z)$, hence they also interlace $\sigma(Z)$ by Lemma˜7.1.
˜(iii) and ˜(iv) are straightforward. ∎
@@
(ii) $\Rightarrow$ (i). Let $X$ be a sprinkled side of $G$. Let $\{P_{1},\ldots,P_{4}\}$ be a partition of the other side V(GX)V(G-X). Let $(A,B)$ be the mixed-4-separation $(P_{1}\cup P_{2}\cup X,X\cup P_{3}\cup P_{4})$ and let $(C,D)$ be the mixed-4-separation $(P_{1}\cup P_{3}\cup X,X\cup P_{2}\cup P_{4})$. Then $(A,B)$ and $(C,D)$ are tetra-separations with separators equal to $X$, and they cross.
(i) $\Rightarrow$ (ii) follows from ˜7.7 ˜(iii) and ˜(iv). ∎